OpenStudy (anonymous):

how do you solve 2x(5 * 4)=3x

OpenStudy (king):

not possible!!!

OpenStudy (anonymous):

2x(20)=3x 40x=3x 40x/3=x Therefore impossible

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

2x(5*4)=3x 2x*20=3x 40x=3x 40x-3x=0 37x=0 x=0